OpenStudy (anonymous):

What is the equation of a line with a slope of –2 that passes through the point (6, 8)?

OpenStudy (mathmale):

Here you're given a point on the line and the slope of the line. Use the point-slope formula for the equation of a straight line: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

Plug in the coordinates of the given point, along with the slope, m. Simplify your result.

OpenStudy (texaschic101):

I use y = mx + b

OpenStudy (texaschic101):

m represents ur slope